# ArchCAD-400K: An Open Large-Scale Architectural CAD Dataset and New Baseline for Panoptic Symbol Spotting

### v2

@@ 5.1 Two-stream Primitive Encoding Module
The extraction of high-quality features for vector graphic panoramic segmentation is crucial as it directly affects segmentation performance. Unlike raster images, vector graphics consist of primitives, yet research on their modeling remains limited, highlighting the need for better feature extraction methods. CADTransformer [12] rasterizes vector graphics and samples features at the primitive centers, while SymPoint [14] encodes primitives as points using geometric attributes and a point cloud encoder[33]. Image-based approaches capture semantics well but struggle with fine-grained tasks, whereas point-based methods handle instances effectively but are less suited for sparse structures like walls. To address this, we combine visual and geometric encoding. Our visual encoder, based on HRNetV2 [34], extracts feature maps from rendered images. Meanwhile, a Point Transformer-inspired point cloud encoder generates geometric features for each primitive, ensuring a more comprehensive representation.
@@ 5.2 Adaptive Fusion Module
where $V_{p}$ is the extracted feature, $\mathcal{K}$ is the interpolation function, and $F_{N(p^{\prime})}$ denotes neighboring pixels. The combination of primitive and image features is challenging due to size variations, noise, and overlapping. To address this, we introduce a geometry-guided fusion layer. First, graphic features $X_{p}$ and image features $V_{p}$ are transformed via an MLP, then concatenated. A weight matrix $w$ is computed as:
@@
where $W_{1}$, $W_{2}$, $W_{3}$ are learnable parameters. Finally, the weighted features yield the final representation:
@@ 5.3 Decoder and Loss Function
The decoder is designed based on the architectures of DETR [35] and Mask2Former [36], achieving panoptic symbol spotting by predicting primitive-level masks along with their categories. The loss function is composed of a standard cross-entropy loss ($L_{cls}$) for class predictions, a binary cross-entropy loss ($L_{bce}$), and a Dice loss ($L_{dice}$[37] for mask predictions. The overall loss is formulated as a weighted sum of the three losses $L=\lambda_{cls}L_{cls}+\lambda_{bce}L_{bce}+\lambda_{dice}L_{dice}$, where $\lambda_{cls}$, $\lambda_{bce}$, and $\lambda_{dice}$ denote the weight for each loss term respectively.
@@ 6 Experiments
We split our ArchCAD-400K dataset into training, validation, and test sets using a 7:1:2 ratio, ensuring that each drawing and its corresponding annotations appear in only one split. This results in 289,144 annotated samples for training, 41,306 for validation, and 82,612 for testing. Following the definition of panoptic symbol spotting, we evaluate the model performance using Panoptic Quality (PQ), Segmentation Quality (SQ), and Recognition Quality (RQ). The formulation of these metrics can be found in [11]. To evaluate model performance on ArchCAD-400K, we compare existing methods and our proposed method DPSS, with detailed results presented below.
@@ 6.1 Implement Details
We adopt HRNetW48 [31] pretrained on COCO-Stuff [38] as the backbone for the visual branch. The point cloud branch uses the PointTransformerV2 [39] encoder pretrained on ScanNetV2 [40]. The model is trained on 8 NVIDIA A800 GPUs and the optimizer is AdamW. On FloorplanCAD [11] dataset, we set a batch size of 2 per GPU and with a learning rate of $2\times 10^{-4}$ and a weight decay of 0.1. The model is trained for 50 epochs. On ArchCAD-400K, we set a batch size of 4 per GPU and train the model for 10 epochs with a learning rate of $2\times 10^{-4}$ and the same weight decay. Other baseline methods on ArchCAD-400K are trained for 10 epochs under their default configurations.
@@ 6.2 Quantitative Evaluation
SymPointV2[13] w/ 90.1 96.3 93.6 90.8 96.6 94.0 80.8 90.9 88.9
CADSpotting[16] w/ 88.9 95.6 93.0 89.7 96.2 93.2 80.6 89.7 89.8
@@
CADTransformer[12] w/o 68.9 88.3 73.3 78.5 94.0 83.5 58.6 81.9 71.5
GAT-CADNet[15] w/o 73.7 91.4 80.7
SymPoint[14] w/o 83.3 91.4 91.1 84.1 94.7 88.8 48.2 69.5 69.4
SymPointV2[13] w/o 83.2 91.3 91.1 85.8 92.5 92.7 49.3 70.3 70.1
@@
Table 2: Panoptic symbol spotting results on FloorplanCAD [11] dataset
@@
SymPointV2 [13] w/ 89.5 88.3 71.3 60.7 60.1
CADSpotting [16] w/ 93.5 93.9 72.2 69.1 69.0
@@
DeepLabv3+R101 [41] w/o 68.8 71.4
DINO [42] w/o 64.9 54.9 47.5
CADTransformer [11] w/o 82.2 80.1
SymPoint [14] w/o 86.8 85.5 66.3 55.7 52.8
SymPointV2 [13] w/o 87.0 86.3 66.4 57.7 57.5
@@
Table 3: Semantic and instance spotting results on FloorPlanCAD
Quantitative comparison on FloorPlanCAD. We evaluate the performance of various methods on the FloorPlanCAD dataset, as shown in Table 2 and Table 3, focusing on both scenarios: one with prior information (such as layers and color) and one without. This dual evaluation is motivated by the intrinsic properties of CAD drawings, where layer and color information naturally encode semantic relationships among graphical primitives. For instance, primitives within the same layer or sharing similar colors are more likely to belong to the same semantic category.
Notably, SymPointV2 extends SymPointV1 by incorporating layer feature encoding, while CADSpotting relies on dense sampling of primitives based on color and positional information. The experimental results in Table 2 show that our DPSS performs similarly to these methods when layer and color information is available. However, in the absence of such priors, SymPointV2 experiences a significant performance drop, while DPSS remains stable, achieving a 3% higher total PQ than the second-best method. Particularly, DPSS outperforms SymPointV2 by 15% in PQ for the "Stuff" category, highlighting its robustness and generalization capabilities. As shown in Table 3, when layer and color information is unavailable, DPSS significantly surpasses the second-best method in both Semantic Spotting and Instance Spotting, further underscoring its superior performance across various metrics.
While our results demonstrate strong performance with the inclusion of prior information, we place greater emphasis on performance without these additional priors. This is because the absence of layer and color data more closely reflects real-world conditions, where such information is not often available.
@@
Method Total Thing Stuff
PQ SQ RQ PQ SQ RQ PQ SQ RQ
CADTransformer[12] 60.0 89.7 66.9 52.5 83.6 62.7 70.1 96.7 72.5
SymPoint[14] 47.6 86.1 55.3 51.4 91.9 55.9 39.9 73.9 54.0
SymPointV2[13] 60.5 88.0 68.8 62.4 91.7 68.1 52.8 73.7 71.7
DPSS 70.6 90.2 78.2 65.6 92.4 70.9 77.6 87.8 88.4
@@
Table 4: Panoptic symbol spotting results on ArchCAD-400K
Method Semantic spotting Instance spotting
F1 wF1 AP50 AP75 mAP
CADTransformer[12] 84.1 83.4
SymPoint[14] 76.8 62.0 36.1 30.9 30.8
SymPointV2[13] 69.8 69.3 44.9 40.2 39.7
DPSS 87.8 84.1 45.6 41.1 40.7
Table 5: Semantic and instance spotting results on ArchCAD-400K
We also evaluate the performance of various methods on our ArchCAD-400K dataset, as shown in Table 4 and Table 5, comparing DPSS with other methods, including CADTransformer, SymPoint, and SymPointV2. Importantly, the results presented here are based on the scenario where layer and color prior information is not utilized, reflecting more realistic conditions.
ArchCAD-400K, being more challenging than FloorPlanCAD, presents additional complexities, which is reflected in the generally lower performance across all methods. For example, SymPointV2, which achieves a total PQ of 83.2% on FloorPlanCAD, shows a significant drop in performance on ArchCAD-400K, with the total PQ dropping by more than 22%, down to 60.5%.
For semantic and instance spotting, DPSS shows significant improvements over other methods, particularly in terms of scalability and robustness. In semantic spotting, DPSS achieves an F1 score of 87.8%, outperforming the CADTransformer (84.1%) and SymPoint (76.8%), and showing a substantial gain of nearly 18% over SymPointV2 (69.8%). In instance spotting, DPSS achieves an mAP of 40.7%, outpacing SymPointV2 (39.7%) and SymPoint (30.8%). For panoptic symbol spotting, DPSS also leads the field, achieving a total PQ of 70.6%, notably higher than SymPointV2 (60.5%) and SymPoint (47.6%). DPSS excels particularly in the Stuff category, where it achieves a Stuff PQ of 77.6%, representing a 24.8% improvement over SymPointV2. These results underscore DPSS’s robustness and ability to generalize effectively, especially in challenging categories that are harder to define.
In summary, even without additional prior information, DPSS demonstrates superior performance across all major metrics compared to other methods, reflecting its strong scalability and robustness. Our ArchCAD-400K dataset, being more challenging, shows that DPSS is not only capable of achieving high performance in more complex scenarios, but also adaptable to real-world applications where such prior information is seldom accessible. These findings position DPSS as a leading method for panoptic symbol spotting in architectural CAD drawings.
6.3 Dataset Ablations
@@
Generalization ability of ArchCAD-400K. To validate the generalization ability of ArchCAD-400K, We implement DPSS on FloorPlanCAD[11] dataset with a dual-pathway encoder pretrained on ArchCAD-400K. Results are illustrated in Figure 8. The pre-trained model achieves faster and more stable convergence, surpassing the performance of non-pre-trained models at 50 epochs within 30 epochs. It suggests that ArchCAD-400K covers FloorPlanCAD and furthermore has more diverse patterns and properties for models to generalize across datasets.
@@
Scaling performance of ArchCAD-400K. We conducted experiments to validate the scaling law in the context of panoptic symbol spotting. Using a subset of our ArchCAD-400K and adopting DPSS as the baseline network, we evaluated the training performance across different data scales, ranging from 10K samples to the full 400K samples. As demonstrated in Figure 8 and 8, within 10K to 400K data range, doubling the dataset size consistently reduces the loss by 2.33 and improves the PQ metric by 6.40% in average. The results indicate that, for panoptic symbol spotting tasks, large-scale datasets significantly enhance model performance.
@@ 6.4 Qualitative Results
The qualitative results on our ArchCAD-400K are illustrated in Figure 9. In the presented cases, DPSS demonstrates superior performance than SymPointV2 [13]. In architectural diagrams (row 1), characterized by a diversity of component types and numerous interfering lines, DPSS is capable of filtering out such disturbances to achieve precise spotting results. In structural drawings (row 2), where the shapes of components exhibit high similarity, DPSS effectively integrates contextual information to discern semantic differences between analogous components. More qualitative results on ArchCAD-400K can be found in our supplementary materials.
(a) GT
(b) DPSS
(c) SymPointV2 [13]
Figure 9: Qualitative comparison between DPSS and SymPointV2 on ArchCAD-400K. (a): Ground truth. (b): Results from DPSS. (c): Results from SymPointV2. Best viewed in color.
@@ 7 Conclusion
Appendix A Detailed description of categorization
In Figure 10, we present some visual examples of content from each category. Engineering symbols from different categories can share great similarity. For example, simple rectangles can represent columns, holes, foundation, or furniture. Similarly, pairs of parallel lines might denote pipeline, beam, or wall. At the same time, some instance can be drawn in diverse forms. As illustrated in Figure 10, there are at least six different ways to represent a door, and the appearance of stairs can vary much in different scenarios.
@@ Appendix A Detailed description of categorization
Appendix B Additional information about the annotation pipeline
@@
Appendix C Additional Examples of ArchCAD-400K
Figure 2 shows an example of well-labeled large architectural drawings. We further show two labeled drawings in our dataset in Figure 11. Each of them covers an area over 8000 m2, with various types of components.
@@ Appendix C Additional Examples of ArchCAD-400K
Appendix D Additional Quantitative Evaluations
We present the detailed experimental results of various methods on the ArchCAD-400K, including the panoptic quality (PQ) for each category as well as the mean IoU (Intersection over Union) for each category. The IoU for each category is obtained by calculating the intersection over union between the predicted panoptic segmentation masks and the ground truth masks. Different methods excel at handling different types of objects, DPSS achieves higher PQ and IoU metrics. For the current spotting results, there is still significant room for improvement.
Table 7: Quantitative results for panoptic symbol spotting of each category on ArchCAD-400K
class DPSS CADTransformer [11] SymPoint [14] SymPointV2 [13]
IoU PQ IoU PQ IoU PQ IoU PQ
symbol 84.4 89.8 62.4 70.7 76.4 80.9 74.2 79.3
axis 90.8 85.4 23.2 32.9 58.3 56.2 71.8 83.7
door 64.4 75.9 53.2 64.9 63.9 75.6 82.6 79.1
floor slab 40.8 85.6 14.4 80.3 20.9 82.8 47.7 69.4
elevator 45.6 77.6 26.5 63.9 54.6 73.3 28.0 69.2
stairs 60.4 65.4 35.3 45.2 41.3 55.4 43.7 78.0
furniture 72.5 88.7 66.5 73.3 63.2 83.9 43.4 54.0
hole 61.1 70.2 32.1 40.0 59.6 53.8 50.2 60.7
window 47.8 57.2 52.4 55.0 64.7 68.4 57.8 70.8
curtain wall 65.9 83.7 48.8 61.4 50.3 72.6 33.9 39.1
wall 73.4 75.6 39.1 45.9 52.7 57.7 42.3 70.7
concrete column 65.8 83.6 59.5 66.8 64.4 71.6 63.9 69.8
steel column 48.6 80.3 50.7 71.8 52.9 82.8 51.1 79.2
concrete beam 81.8 81.7 43.8 45.0 61.5 62.1 48.3 77.9
steel beam 79.1 73.0 19.7 34.9 47.5 52.2 71.5 77.4
parking spot 66.8 73.2 57.9 76.9 46.0 79.3 72.1 69.0
roof 3.5 28.7 3.7 39.1 5.7 50.1 55.6 76.1
base 85.0 90.3 65.8 75.2 71.2 83.1 16.3 21.3
bracing 19.6 33.8 7.8 20.0 20.2 23.2 67.6 81.1
rebar 72.7 93.1 39.4 69.4 56.5 81.1 53.5 75.9
equipment 37.2 45.8 1.0 7.0 30.8 31.8 28.4 29.1
handrail 73.6 65.9 31.2 31.0 40.2 42.7 70.3 85.8
pipe 44.4 48.3 32.9 34.4 51.7 41.7 52.5 42.4
window curtain 50.4 80.3 58.6 72.1 49.2 67.7 46.3 47.5
construction joint 0.0 5.3 0.0 0.3 2.5 4.8 4.1 3.7
embedment 0.0 0.0 0.0 0.0 0.0 0.0 1.2 2.8
hydrant 70.7 71.7 68.6 64.5 74.1 69.4 49.5 34.0
overall 70.6 67.04 47.6 49.1 60.5 59.1 60.0 60.8
Appendix E Additional Qualitative Evaluations
Qualitative results on FloorPlanCAD [11] are shown in Figure 12, DPSS performs better despite the noise of the filler graphics and overlaps between instances.
(a) GT
(b) DPSS
(c) SymPointV2[13]
(d) GT
(e) DPSS
(f) SymPointV2[13]
Figure 12: Additional qualitative comparison between DPSS and SymPointV2 on ArchCAD-400K. (a) and (d): Panoptic annotations. (b) and (e): Results from DPSS. (c) and (f): Results from SymPointV2.

### v3

@@ 5.1 Two-stream Primitive Encoding Module
The extraction of high-quality features for vector graphic panoramic segmentation is crucial as it directly affects segmentation performance. Unlike raster images, vector graphics consist of primitives, yet research on their modeling remains limited, highlighting the need for better feature extraction methods. CADTransformer fan2022cadtransformer rasterizes vector graphics and samples features at the primitive centers, while SymPoint liu2024symbol encodes primitives as points using geometric attributes and a point cloud encoderzhao2021point. Image-based approaches capture semantics well but struggle with fine-grained tasks, whereas point-based methods handle instances effectively but are less suited for sparse structures like walls. To address this, we combine visual and geometric encoding. Our visual encoder, based on HRNetV2 wang2020deep, extracts feature maps from rendered images. Meanwhile, a Point Transformer-inspired point cloud encoder generates geometric features for each primitive, ensuring a more comprehensive representation.
@@ 5.2 Adaptive Fusion Module
where $V_{p}$ is the extracted feature, $\mathcal{K}$ is the interpolation function, and $F_{N(p^{\prime})}$ denotes neighboring pixels. The combination of primitive and image features is challenging due to size variations, noise, and overlapping. To address this, we introduce a geometry-guided fusion layer. First, graphic features $X_{p}$ and image features $V_{p}$ are transformed via an MLP, then concatenated and pass through an activation function $\sigma$. A weight matrix $w$ is computed as:
@@
where $W_{1}$, $W_{2}$, $W_{3}$ are learnable parameters. Finally, the weighted features yield the representation:
@@ 5.3 Decoder and Loss Function
The decoder is designed based on the architectures of DETR carion2020end and Mask2Former cheng2022masked, achieving panoptic symbol spotting by predicting primitive-level masks along with their categories. The loss function is composed of a standard cross-entropy loss ($L_{cls}$) for class predictions, a binary cross-entropy loss ($L_{bce}$), and a Dice loss ($L_{dice}$milletari2016v for mask predictions. The overall loss is formulated as a weighted sum of the three losses $L=\lambda_{cls}L_{cls}+\lambda_{bce}L_{bce}+\lambda_{dice}L_{dice}$, where $\lambda_{cls}$, $\lambda_{bce}$, and $\lambda_{dice}$ denote the weight for each loss term respectively.
(a) DPSS Framework
(b) Adaptive fusion
Figure 5: The overall architecture of the proposed Dual-Pathway Symbol Spotter (DPSS). CAD drawings are encoded through both point backbone and image backbone. Features from dual pathways are aligned and fused in adaptive fusion module, then pass through a transformer decoder to generate mask predictions and classifications at primitive level.
@@ 6 Experiments
We split our ArchCAD-400k dataset into training, validation, and test sets using a 7:1:2 ratio, ensuring that each drawing and its corresponding annotations appear in only one split. This results in 289,144 annotated samples for training, 41,306 for validation, and 82,612 for testing. Following the definition of panoptic symbol spotting, we evaluate the model performance using Panoptic Quality (PQ), Segmentation Quality (SQ), and Recognition Quality (RQ). The formulation of these metrics can be found in fan2021floorplancad. To evaluate model performance on ArchCAD-400k, we compare existing methods and our proposed method DPSS, with detailed results presented below.
@@ 6.1 Implement Details
We adopt HRNetW48 sun2019deep pretrained on COCO-Stuff caesar2018coco as the backbone for the visual branch. The point cloud branch uses the PointTransformerV2 wu2022point encoder pretrained on ScanNetV2 dai2017scannet. The model is trained on 8 NVIDIA A800 GPUs and the optimizer is AdamW. On FloorplanCAD fan2021floorplancad dataset, we set a batch size of 2 per GPU and with a learning rate of $2\times 10^{-4}$ and a weight decay of 0.1. The model is trained for 50 epochs. On ArchCAD-400k, we set a batch size of 4 per GPU and train the model for 10 epochs with a learning rate of $2\times 10^{-4}$ and the same weight decay. Other baseline methods on ArchCAD-400k are trained for 10 epochs under their default configurations. All models were confirmed to have converged before evaluation to ensure fair performance comparison.
@@ 6.2 Quantitative Evaluation
Table 2: Panoptic symbol spotting results on FloorplanCAD fan2021floorplancad dataset
@@
SymPointV2liu2024sympoint w/ 90.1 96.3 93.6 90.8 96.6 94.0 80.8 90.9 88.9
CADSpottingmu2024cadspotting w/ 88.9 95.6 93.0 89.7 96.2 93.2 80.6 89.7 89.8
@@
CADTransformerfan2022cadtransformer w/o 68.9 88.3 73.3 78.5 94.0 83.5 58.6 81.9 71.5
GAT-CADNetzheng2022gat w/o 73.7 91.4 80.7
SymPointliu2024symbol w/o 83.3 91.4 91.1 84.1 94.7 88.8 48.2 69.5 69.4
SymPointV2liu2024sympoint w/o 83.2 91.3 91.1 85.8 92.5 92.7 49.3 70.3 70.1
@@
Table 3: Semantic and instance spotting results on FloorPlanCAD
@@
SymPointV2 liu2024sympoint w/ 89.5 88.3 71.3 60.7 60.1
CADSpotting mu2024cadspotting w/ 93.5 93.9 72.2 69.1 69.0
@@
DeepLabv3+R101 chen2017rethinking w/o 68.8 71.4
DINO zhang2022dino w/o 64.9 54.9 47.5
CADTransformer fan2021floorplancad w/o 82.2 80.1
SymPoint liu2024symbol w/o 86.8 85.5 66.3 55.7 52.8
SymPointV2 liu2024sympoint w/o 87.0 86.3 66.4 57.7 57.5
@@
Quantitative comparison on FloorPlanCAD. We evaluate multiple methods on the FloorPlanCAD dataset (Tables 2 and 3), under two scenarios: with and without prior information (such as layers and color). This distinction reflects how CAD drawings encode semantics—primitives in the same layer or with similar colors often share categories.
SymPointV2 improves on SymPointV1 by adding layer encoding, while CADSpotting samples primitives based on color and position. As shown in Table 2, DPSS performs comparably to these methods when priors are present. Without them, however, SymPointV2 suffers a notable drop, whereas DPSS stays robust—achieving 3% higher overall PQ and 15% higher "Stuff" PQ than SymPointV2, demonstrating better generalization. In Table 3, DPSS also clearly outperforms others in Semantic and Instance Spotting without priors, including some classical semantic segmentation and instance detection algorithms.
While our method benefits from prior information, we prioritize performance without it, as this better reflects real-world scenarios where such metadata is often unavailable.
@@
Table 4: Panoptic, semantic, and instance symbol spotting results on ArchCAD-400k
Method Total Thing Stuff Semantic Spotting Instance Spotting
PQ SQ RQ PQ SQ RQ PQ SQ RQ F1 wF1 AP50 AP75 mAP
CADTransformerfan2022cadtransformer 60.0 89.7 66.9 52.5 83.6 62.7 70.1 96.7 72.5 84.1 83.4
SymPointliu2024symbol 47.6 86.1 55.3 51.4 91.9 55.9 39.9 73.9 54.0 76.8 62.0 36.1 30.9 30.8
SymPointV2liu2024sympoint 60.5 88.0 68.8 62.4 91.7 68.1 52.8 73.7 71.7 69.8 69.3 44.9 40.2 39.7
DPSS 70.6 90.2 78.2 65.6 92.4 70.9 77.6 87.8 88.4 87.8 84.1 45.6 41.1 40.7
We evaluate various methods on our ArchCAD-400k dataset (Table 4), comparing DPSS with CADTransformer, SymPoint, and SymPointV2, all without using layer or color priors, to better reflect real-world conditions. Compared to FloorPlanCAD, ArchCAD-400k is more challenging, leading to overall lower performance. For instance, SymPointV2 drops from 83.2% PQ on FloorPlanCAD to 60.5% on ArchCAD-400k —a decrease of over 22%.
DPSS shows clear advantages in scalability and robustness. In semantic spotting, it achieves an F1 of 87.8%, outperforming CADTransformer (84.1%), SymPoint (76.8%), and SymPointV2 (69.8%). In instance spotting, it reaches 40.7% mAP, surpassing SymPointV2 (39.7%) and SymPoint (30.8%). For panoptic symbol spotting, DPSS attains a total PQ of 70.6%, significantly ahead of SymPointV2 (60.5%) and SymPoint (47.6%). Notably, its Stuff PQ reaches 77.6%, 24.8% higher than SymPointV2.
In summary, DPSS consistently outperforms baselines across all tasks, even without prior information. Its strong performance on the more complex ArchCAD-400k highlights its scalability, robustness, and practical applicability to real-world CAD scenarios.
6.3 Ablations
Ablation studies on DPSS. To demonstrate the effectiveness of the proposed DPSS, we conducted ablation studies focusing on different encoding strategies. Specifically, we evaluated the impact of using only the primitive encoder, only the image encoder, and the combination of both, as shown in Table 5 (Lines 1–3). The integration of both the image encoder and the primitive encoder yields a 1.3% improvement in PQ compared to using the primitive encoder alone, and a 2.1% gain over using only the image encoder. These results highlight the complementary nature of the two encoding branches. Furthermore, we investigated the role of the adaptive fusion module by replacing it with a simple concatenation strategy for combining image and primitive features, as shown in Table 5 (Lines 3–4). The results show that incorporating our adaptive fusion leads to a significant 3.2% increase in PQ, underscoring its effectiveness in enhancing feature integration.
Table 5: Ablation experiments on FloorPlanCAD
Primitive Encoder Image Encoder Adaptive Fusion PQ RQ SQ
81.7 90.1 90.6
80.9 89.4 90.6
83.0 90.1 92.1
86.2 93.0 92.6
@@ 6.3 Ablations
Generalization ability of ArchCAD-400k. To validate the generalization ability of ArchCAD-400k, We implement DPSS on FloorPlanCADfan2021floorplancad dataset with a dual-pathway encoder pretrained on ArchCAD-400k. Results are illustrated in Figure 6. The pre-trained model achieves faster and more stable convergence, surpassing the performance of non-pre-trained models at 50 epochs within 30 epochs. It suggests that ArchCAD-400k covers FloorPlanCAD and furthermore has more diverse patterns and properties for models to generalize across datasets.
@@
Scaling performance of ArchCAD-400k. We conducted experiments to validate the scaling law in the context of panoptic symbol spotting. Using a subset of our ArchCAD-400k and adopting DPSS as the baseline network, we evaluated the training performance across different data scales, ranging from 10K samples to the full 400K samples. As demonstrated in Figure 6 and 6, within 10K to 400K data range, doubling the dataset size consistently reduces the loss by 2.33 and improves the PQ metric by 6.40% in average. The results indicate that, for panoptic symbol spotting tasks, large-scale datasets significantly enhance model performance.
@@
(a) Ground truth
(b) DPSS
(c) SymPointV2
(d) Ground truth
(e) DPSS
(f) SymPointV2
Figure 7: Qualitative comparison between DPSS and SymPointV2 liu2024sympoint on FloorPlanCAD(Line1) and ArchCAD-400k (Line2).
@@ 6.4 Qualitative Results
The qualitative results on our ArchCAD-400k are illustrated in Figure 7. In the presented cases, DPSS demonstrates superior performance than SymPointV2 liu2024sympoint. In architectural diagrams (row 1), characterized by a diversity of component types and numerous interfering lines, DPSS is capable of filtering out such disturbances to achieve precise spotting results. In structural drawings (row 2), where the shapes of components exhibit high similarity, DPSS effectively integrates contextual information to discern semantic differences between analogous components.
@@ 7 Conclusion
Appendix A More Information About the ArchCAD-400k
A.1 Detailed description of categorization
In Figure 8, we present some visual examples of content from each category. Engineering symbols from different categories can share great similarity. For example, simple rectangles can represent columns, holes, foundation, or furniture. Similarly, pairs of parallel lines might denote a pipeline, beam, or wall. At the same time, some instances can be drawn in diverse forms. As illustrated in Figure 8, there are at least six different ways to represent a door, and the appearance of stairs can vary greatly in different scenarios.
@@ A.1 Detailed description of categorization
A.2 Additional information about the annotation pipeline
@@
A.3 Additional examples of ArchCAD-400k
An example of well-labeled large architectural drawings exists in the main text. We further show two labeled drawings in our dataset in Figure 9. Each of them covers an area over 8000 m2, with various types of components.
@@ A.3 Additional examples of ArchCAD-400k
Appendix B More Evaluations
B.1 Analysis on Computational Efficiency
To evaluate the computational efficiency of the proposed method, we conducted additional experiments measuring inference latency across representative approaches. Our model operates on images with a resolution of 700×700, which introduces only a moderate computational overhead. Compared with SymPoint-V2 [liu2024sympoint], the proposed method increases inference latency by approximately 30%, while achieving 3.0% and 10.1% higher panoptic quality on ArchCAD-400K and FloorPlanCAD, respectively. These results indicate that the proposed approach attains a favorable balance between computational efficiency and segmentation accuracy.
Table 7: Comparison with existing methods on FloorPlanCAD and ArchCAD-400K.
Method PQ (FloorPlanCAD) PQ (ArchCAD-400K) Average Latency
CADTransformer [fan2022cadtransformer] 68.9 60.0 159ms
SymPoint [liu2024symbol] 83.3 47.6 66ms
SymPoint-V2 [liu2024sympoint] 83.2 60.5 95ms
DPSS (ours) 86.2 70.6 121ms
B.2 Baseline Results on the Public Subset of ArchCAD-400k
To facilitate further research in this field, we release a publicly available subset of the ArchCAD-400k dataset, which has been further refined and optimized from the original version. While this open-source release preserves the overall structure of the initial dataset, minor differences exist in the annotation details due to enhanced preprocessing and quality control. We also train several representative models on this subset and report the corresponding baseline results for reference.
Table 8: Evaluation results of multiple methods on the publicly available ArchCAD-400k.
Method Total Thing Stuff
PQ SQ RQ PQ SQ RQ PQ SQ RQ
CADTransformer [fan2021floorplancad] 64.3 72.3 88.8 66.5 72.4 92.9 56.6 72.1 78.5
SymPoint  [liu2024symbol] 67.7 75.3 90.0 72.3 78.4 92.2 43.0 58.2 73.8
SymPointV2 [liu2024sympoint] 76.6 84.6 90.6 80.0 85.6 93.1 61.0 79.2 77.1
DPSS (ours) 82.9 93.3 88.8 86.3 93.6 92.2 78.4 93.1 84.2
Appendix C Limitations and Future Work
Although the proposed DPSS demonstrates strong performance on the panoptic symbol spotting task, some limitations remain. Notably, the current approach cannot process an entire vector drawing in a single pass, which leads to high computational overhead. With the availability of the ArchCAD-400k dataset, more complex and comprehensive research questions related to panoptic symbol spotting and the analysis of engineering line drawings can be explored. For example, future directions include the pre-trained models tailored for CAD vector graphics, and efficient inference strategies to handle large scale real-world drawings.